High Desert Potteryworks makes a variety of pottery products that it sells to retailers such as Home Depot. The company uses a job-order costing system in which predetermined overhead rates are used to apply manufacturing overhead cost to jobs. The predetermined overhead rate in the Molding Department is based on machine-hours, and the rate in the Painting Department is based on direct labor-hours. At the beginning of the year, the company's management made the following estimates:

Department
   Molding Painting
  Direct labor-hours 12,000    60,000   
  Machine-hours 70,000    8,000   
  Direct materials cost $510,000    $650,000   
  Direct labor cost $130,000    $420,000   
  Fixed manufacturing overhead cost $497,000    $615,000   
  Variable manufacturing overhead per machine-hour $1.50    -      
  Variable manufacturing overhead per direct labor-hour -       $2.00   

   Job 205 was started on August 1 and completed on August 10. The company's cost records
show the following information concerning the job:

Department
   Molding Painting
  Direct labor-hours 30      84     
  Machine-hours 110      20     
  Materials placed into production $470      $332     
  Direct labor cost $325      $588     

   Required:

1.

Compute the predetermined overhead rate used during the year in the Molding Department. Compute the rate used in the Painting Department. (Round your answers to 2 decimal places.)

Predetermined overhead rate
Molding Department per MH
Painting Department per DLH

Compute the total overhead cost applied to Job 205. (Round "Predetermined overhead rate" to 2 decimal places, other intermediate calculations and final answers to the nearest dollar amount.)

Total overhead cost
         

What would be the total cost recorded for Job 205? (Round "Predetermined overhead rate" to 2 decimal places, other intermediate calculations and final answers to the nearest dollar amount.)

Department
Molding Painting Total
Direct materials
Direct labor
Manufacturing overhead
Total cost

If the job contained 50 units, what would be the unit product cost? (Round "Predetermined overhead rate" and final answer to 2 decimal places and other intermediate calculations to the nearest dollar amount.)

Unit product cost           per unit

At the end of the year, the records of High Desert Potteryworks revealed the following actual cost and operating data for all jobs worked on during the year:

Department
   Molding Painting
  Direct labor-hours 10,000      62,000     
  Machine-hours 65,000      9,000     
  Direct materials cost $430,000      $680,000     
  Direct labor cost $108,000      $436,000     
  Manufacturing overhead cost $570,000      $750,000     

What was the amount of underapplied or overapplied overhead in each department at the end of the year? (Round "Predetermined overhead rate" to 2 decimal places, other intermediate calculations and final answers to the nearest dollar amount. Select "None" for no effect (i.e., zero variance).)

Molding Department                          
Painting Department
Answer #1
Concepts and reason

Job order costing: It is a method of cost accounting, in which cost is collected and accumulated for each job, work order, or project separately. Especially the job order costing is followed in organizations where customized goods are produced.

Variable manufacturing overhead: Variable manufacturing overhead is a type of manufacturing overhead, which are changed according to the level of production.

Fixed manufacturing overhead: It is a type of manufacturing overhead; it doesn’t change according to the level of production.

Fundamentals

Manufacturing overhead costs: The costs, which do not relate directly with the manufacturing of products, are referred to as manufacturing overhead costs or indirect costs.

Predetermined overhead rate: The cost accountants estimate a single overhead rate for all the overhead costs to compute manufacturing overhead. That allocation rate, which is computed as a ratio of total estimated overhead costs to total estimated quantity, is known as predetermined overhead rate.

Unit product cost: Total manufacturing cost divided by number of units produced is the unit product cost for the job.

Total costs: The total costs that are incurred to manufacture goods are referred to as total costs. Total cost is the sum of direct materials, direct labor and manufacturing overhead.

Direct labor cost: The costs that are related to the labor employed in manufacturing process are known as direct labor costs. The direct labor cost includes not only wages of employees, but also worker’s compensation.

Direct materials: Direct materials are the raw materials which are directly related with the production of the goods.

Underapplied overhead: Excess of actual manufacturing overhead costs over the manufacturing overhead applied is the underapplied overhead of the company.

Overapplied overhead: Excess of manufacturing overhead applied over the manufacturing overhead incurred is the overapplied overhead of the company.

1)

Calculate predetermined overhead rate.

Predetermined overhead rate
Molding Department
Painting Department
S 8.60 per MH
S 12.25 per DLH

Working note:

Calculate predetermined overhead rate of moulding department.

Predeterminedoverheadrate=((Estimatedfixedmanufacturingoverhead+(Machinehours×(Variablemanufacturingoverheadpermachinehour)))Estimatedtotalmachinehours)=(($497,000+(70,000machinehours×$1.5machinehour))70,000machinehours)=$602,00070,000machinehours=$8.60permachinehour\begin{array}{c}\\{\rm{Predetermined overhead rate = }}\left( {\frac{{\left( \begin{array}{l}\\{\rm{Estimated fixed manufacturing overhead }}\\\\{\rm{ + }}\left( \begin{array}{l}\\{\rm{Machine hours }}\\\\ \times {\rm{ }}\left( \begin{array}{l}\\{\rm{Variable manufacturing overhead}}\\\\{\rm{ per machine hour}}\\\end{array} \right)\\\end{array} \right)\\\end{array} \right)}}{{{\rm{Estimated total machine hours}}}}} \right){\rm{ }}\\\\{\rm{ = }}\left( {\frac{{\left( {{\rm{\$ 497,000 + }}\left( \begin{array}{l}\\{\rm{70,000 machine hours }}\\\\ \times \;\$ {\rm{1}}{\rm{.5 machine hour}}\\\end{array} \right)} \right)}}{{{\rm{70,000 machine hours}}}}} \right){\rm{ }}\\\\{\rm{ = }}\frac{{{\rm{\$ 602,000}}}}{{{\rm{70,000 machine hours}}}}{\rm{ }}\\\\{\rm{ = \$ 8}}{\rm{.60 per machine hour}}\\\end{array}

Calculate predetermined overhead rate of painting department.

Predeterminedoverheadrate=((Estimatedfixedmanufacturingoverhead+(DLHs×(VariablemanufacturingoverheadperDLHs)))EstimatedtotalDLHs)=(($615,000+(60,000DLHs×$2DLH))60,000DLHs)=$735,00060,000DLHs=$12.25perDLH\begin{array}{c}\\{\rm{Predetermined overhead rate = }}\left( {\frac{{\left( \begin{array}{l}\\{\rm{Estimated fixed manufacturing overhead }}\\\\{\rm{ + }}\left( \begin{array}{l}\\{\rm{DLHs }}\\\\{\rm{ \times }}\left( \begin{array}{l}\\{\rm{Variable manufacturing overhead}}\\\\{\rm{ per DLHs}}\\\end{array} \right)\\\end{array} \right)\\\end{array} \right)}}{{{\rm{Estimated total DLHs}}}}} \right){\rm{ }}\\\\{\rm{ = }}\left( {\frac{{\left( {{\rm{\$ 615,000 + }}\left( \begin{array}{l}\\{\rm{60,000 DLHs }}\\\\{\rm{ \times }}\;{\rm{\$ 2 DLH}}\\\end{array} \right)} \right)}}{{{\rm{60,000 DLHs}}}}} \right){\rm{ }}\\\\{\rm{ = }}\frac{{{\rm{\$ 735,000}}}}{{{\rm{60,000 DLHs}}}}{\rm{ }}\\\\{\rm{ = \$ 12}}{\rm{.25 per DLH}}\\\end{array}

Calculate total overhead cost applied to Job 205.

TotaloverheadcostappliedtoJob205}=(Moldingdepartmentoverheadapplied+Paintingdepartmentoverheadapplied)=$946+$1,029=$1,975\begin{array}{c}\\\left. \begin{array}{l}\\{\rm{Total overhead cost}}\\\\{\rm{ applied to Job 205}}\\\end{array} \right\}{\rm{ = }}\left( \begin{array}{l}\\{\rm{Molding department overhead applied }}\\\\{\rm{ + Painting department overhead applied}}\\\end{array} \right){\rm{ }}\\\\{\rm{ = \$ 946 + \$ 1,029 }}\\\\{\rm{ = \$ 1,975}}\\\end{array}

Therefore, the total overhead cost applied to Job 205 is $1,975.

Working note:

Calculate moulding department overhead applied.

Moldingdepartmentoverheadapplied=(Machinehours×Predeterminedoverheadrate)=110MHs×$8.60perMH=$946\begin{array}{c}\\{\rm{Molding department overhead applied = }}\left( \begin{array}{l}\\{\rm{Machine hours }}\\\\ \times {\rm{ Predetermined overhead rate}}\\\end{array} \right){\rm{ }}\\\\{\rm{ = 110 MHs }} \times {\rm{ \$ 8}}{\rm{.60 per MH }}\\\\{\rm{ = \$ 946}}\\\end{array}

Calculate painting department overhead applied.

Paintingdepartmentoverheadapplied=(Directlaborhours×Predeterminedoverheadrate)=84DLHs×$12.25perDLH=$1,029\begin{array}{c}\\{\rm{Painting department overhead applied = }}\left( \begin{array}{l}\\{\rm{Direct labor hours }}\\\\ \times {\rm{ Predetermined overhead rate}}\\\end{array} \right){\rm{ }}\\\\{\rm{ = 84 DLHs }} \times {\rm{ \$ 12}}{\rm{.25 per DLH }}\\\\{\rm{ = \$ 1,029}}\\\end{array}

Calculate total cost recorded for Job 205.

Molding PaintingTotal
$ 332
S 588
$1,029 1,975
S 1,741 $1,949|3,690
Direct materials
Direct labor
Manufacturing overhead
Tota

Therefore, the total cost recorded for Job 205 is $3,690.

Calculate unit product cost.

Unitproductcost=TotalmanufacturingcostNumberofunitsinthejob=$3,69050units=$73.80perunit\begin{array}{c}\\{\rm{Unit product cost = }}\frac{{{\rm{Total manufacturing cost}}}}{{{\rm{Number of units in the job}}}}{\rm{ }}\\\\{\rm{ = }}\frac{{{\rm{\$ 3,690}}}}{{{\rm{50 units}}}}{\rm{ }}\\\\{\rm{ = \$ 73}}{\rm{.80 per unit}}\\\end{array}

Therefore, the unit product cost is $73.80 per unit.

Calculate amount of underapplied or overapplied overhead in moulding and painting department.

Molding Department
Painting Department
|Underapplied overheadS 11,000|
Overapplied overhead
$ 9,500

Working note:

Calculate underapplied overhead in the moulding department.

Underappliedoverheadinthemoldingdepartment}=(Manufacturingoverheadcostincurred(Machinehours×Predeterminedoverheadrate))=($570,000(65,000MHs×$8.60perMH))=$570,000$559,000=$11,000\begin{array}{c}\\\left. \begin{array}{l}\\{\rm{Underapplied overhead }}\\\\{\rm{in the molding department}}\\\end{array} \right\}{\rm{ = }}\left( \begin{array}{l}\\{\rm{Manufacturing overhead cost incurred }}\\\\ - {\rm{ }}\left( \begin{array}{l}\\{\rm{Machine hours }}\\\\ \times {\rm{ Predetermined overhead rate}}\\\end{array} \right)\\\end{array} \right){\rm{ }}\\\\{\rm{ = }}\left( {{\rm{\$ 570,000 }} - {\rm{ }}\left( \begin{array}{l}\\{\rm{65,000 MHs }}\\\\ \times {\rm{ \$ 8}}{\rm{.60 per MH}}\\\end{array} \right)} \right){\rm{ }}\\\\{\rm{ = \$ 570,000 - \$ 559,000 }}\\\\{\rm{ = \$ 11,000}}\\\end{array}

Therefore, the underapplied overhead is $11,000.

Calculate overapplied overhead in the painting department.

Overappliedoverheadinpaintingdepartment}=((Directlaborhours×Predeterminedoverheadrate)Manufacturingoverheadcostincurred)=((62,000DLHs×$12.25perDLH)$750,000)=$759,500$750,000=$9,500\begin{array}{c}\\\left. \begin{array}{l}\\{\rm{Overapplied overhead }}\\\\{\rm{in painting department}}\\\end{array} \right\}{\rm{ = }}\left( \begin{array}{l}\\\left( \begin{array}{l}\\{\rm{Direct labor hours }}\\\\ \times {\rm{ Predetermined overhead rate}}\\\end{array} \right){\rm{ }}\\\\ - {\rm{ Manufacturing overhead cost incurred}}\\\end{array} \right){\rm{ }}\\\\{\rm{ = }}\left( {\left( \begin{array}{l}\\{\rm{62,000 DLHs }}\\\\ \times {\rm{ \$ 12}}{\rm{.25 per DLH}}\\\end{array} \right){\rm{ }} - {\rm{ \$ 750,000}}} \right){\rm{ }}\\\\{\rm{ = \$ 759,500 }} - {\rm{ \$ 750,000 }}\\\\{\rm{ = \$ 9,500}}\\\end{array}

Therefore, the overapplied overhead is $9,500.

Ans: Part 1

Predetermined overhead rate
Molding Department
Painting Department
S 8.60 per MH
S 12.25 per DLH

Total overhead cost applied to Job 205 is $1,975.

Total cost recorded for Job 205 is $3,690.

Unit product cost is $73.80 per unit.

Molding Department
Painting Department
|Underapplied overheadS 11,000|
Overapplied overhead
$ 9,500